### 14 Mar 2024 v1

@@ 4.1 Experiment Setup
We compare our methods with several baselines for data selection: (1) Random Selection selects data randomly; (2) Instruction Length utilizes length of instruction to determine the importance of an instruction sample; (3) Perplexity computes the perplexity score of an instruction sample to measure its importance; (4) GradN [24] measures the importance of each sample by the L2-norm of the gradient caused by each sample; (5) E2LN [24] measures the importance of each sample by the L2-norm of the error vector of each sample. The E2LN scores are primarily used for estimating sample importance in image classification tasks. To adapt it for visual instruction tuning, we compute all the error vectors for each token in each sample, and then compute the final E2LN score by averaging norms of all error vectors.
@@ 4.2 Main Results
For the traditional data selection approaches used in single-modal classification tasks (GradN and E2LN), the performance is not ideal across most of the benchmarks. A possible reason is that these approaches tend to select samples with high gradient norm. The= selected samples may contain a significant amount of noise or deviate greatly from the model’s optimization direction, which can have a side effect for the model’s performance.

### 21 Mar 2024 v2

@@ 4.1 Experiment Setup
We compare our methods with several baselines for data selection: (1) Random Selection selects data randomly; (2) Instruction Length utilizes length of instruction to determine the importance of an instruction sample; (3) Perplexity computes the perplexity score of an instruction sample to measure its importance; (4) GraNd [24] measures the importance of each sample by the L2-norm of the gradient caused by each sample; (5) EL2N [24] measures the importance of each sample by the L2-norm of the error vector of each sample. The EL2N scores are primarily used for estimating sample importance in image classification tasks. To adapt it for visual instruction tuning, we compute all the error vectors for each token in each sample, and then compute the final EL2N score by averaging norms of all error vectors.
@@ 4.2 Main Results
For the traditional data selection approaches used in single-modal classification tasks (GraNd and EL2N), the performance is not ideal across most of the benchmarks. A possible reason is that these approaches tend to select samples with high gradient norm. The= selected samples may contain a significant amount of noise or deviate greatly from the model’s optimization direction, which can have a side effect for the model’s performance.